# LeTac-MPC: Learning Model Predictive Control for Tactile-reactive Grasping

### v1

@@ IV-B Model Predictive Control
TABLE IV: Results of daily objects grasping and transportation experiment. If an experiment fails, we indicate the corresponding table position with red font to highlight the failed result. If the experiment succeeds, we write the average force of grasping measured by the WSG 50-110 gripper in the table position. The bolded font represents the minimum grasping force for the same object. We report the average value and standard deviation of the force during the experimental process after the grasping stabilizes.
@@
Figure 13: Position trajectories for the dynamic transportation (Section V-A), dynamic shaking (Section V-C), and obstacle collision (Section V-D) experiments. The maximum velocity of the trajectory for dynamic transportation is 1.40 m/s, with a maximum acceleration of 3.82 m/$\text{s}^{2}$. For dynamic shaking, the trajectory has a maximum velocity of 1.38 m/s and a maximum acceleration of 8.15 m/$\text{s}^{2}$. Lastly, the trajectory for obstacle collision has a maximum velocity of 0.47 m/s and a maximum acceleration of 0.65 m/$\text{s}^{2}$.
@@
Figure 15: Gripper width curves of dynamic shaking experiments with MPC and PD control.
@@
where $K_{c}$ is a scalar factor. Equation (48) assumes a linear relationship between the contact area $c$ and the position of the gripper $p$, which is a simple and valid approximation of complex gel dynamics. The linear model is a local approximation for a very short duration. In MPC, we don’t using the linear model for long-term predictions, as this would lead to significant errors. Our proposed MPC baseline method runs at a high frequency with receding horizon control, which ensures that we can continuously iterate with the local model to minimize model errors. Combing equations (5) and (48), we can get following model:
@@
$J(\mathbf{y}_{n},\mathbf{a}_{n})$ $=P\mathbf{e}_{n+N}^{T}{\mathbf{Q}}\mathbf{e}_{n+N}+\sum_{k=n}^{n+N-1}\mathbf{e}_{k}^{T}{\mathbf{Q}}\mathbf{e}_{k}+Q_{a}a_{n+m}^{2},$ (62)
@@
The optimization problem (68) is a QP. Unlike the differentiable MPC layer, we don’t need to do backpropagation for the optimization problem (68). Here, we only need to solve the optimization problem (68) to compute the control inputs. Therefore, we use OSQP [48], a efficient solver for QP to solve the optimization problem (68) in real-time. The parameters we use for MPC are shown in Tables II and III.
@@ IV-C Open-loop Grasping
For open-loop grasping, we use the force feedback of the WSG 50-110 gripper as the signal to choose gripper width. We select 10 N as the threshold value. At the beginning of the grasping, the gripper width decreases and once the force feedback is greater than 10 N, the gripper stops. Then the gripper width will remain on this position in the period of performing the task.
@@
It’s important to note that even when we use 10 N as the static condition for the fingers in open-loop grasping, it does not imply that open-loop grasping can maintain a constant grasping force of 10N. This variability is partly due to the inherent noise in the force feedback of the WSG gripper. Furthermore, when the force feedback reaches 10 N, the fingers are in motion, and the force exerted differs once they stops. Additionally, the final resting force varies with the shape and properties of the grasped object, as shown in Table IV.
@@ V Experiments
We conduct experimental validation of proposed LeTac-MPC on a WSG 50-110 parallel gripper and a Franka Panda manipulator, as shown in Fig. LABEL:fig:firstPage. We run LaTac-MPC at a frequency of 25 Hz. Using the official WSG 50 ROS package [^1] https://github.com/nalt/wsg50-ros-pkg, the maximum rate of the gripper is 30 Hz. Therefore, the gripper can track the motion generated by LeTac-MPC with reasonable saturation constraints. We perform three tasks, namely daily objects grasping and transportation (Section V-A), dynamic shaking (Section V-C), and obstacle collision (Section V-D), and compare LeTac-MPC with PD control, MPC, and open-loop grasping. To capture the tactile feedback, we mount a GelSight on one of the fingers of the WSG 50-110 gripper. A video includes these three tasks and comparisons with the baseline methods can be found in the supplemental material.
The WSG gripper we use has a built-in low-level controller, which can track position reference. Our designed reactive grasping controllers are high-level controllers. These high-level controllers produce reference motions for the gripper fingers with reasonable saturation constraints, which are then accurately tracked by the low-level controller. In the case of LeTac-MPC and MPC, the solution to the optimization problem is acceleration $a_{n}$, but our state also includes gripper width $p_{n}$ and velocity $v_{n}$. Eventually, we use $p_{n}$ from the state as the reference sent to the low-level controller for tracking. Although we do not directly track $v_{n}$ and $a_{n}$, since $p$ is derived from the integration of $v$ and $a$, what the low-level controller effectively tracks is the entire trajectory of the finger motion. As for the PD controller, at each moment, we calculate $v_{n}$ through feedback, which, after integration, gives us the position reference $p_{n}$ to be tracked by the low-level controller. Hence, the low-level controller is actually tracking both $p$ and $v$.
@@ V-A Daily Objects Grasping and Transportation
We evaluate our proposed LeTac-MPC method and compare it with three baseline methods for a daily objects grasping and transportation task. As shown in Fig. 6, we select 10 daily objects. These objects differ in physical properties (such as stiffness, total weight, mass distribution, and coefficient of friction), sizes and shapes (ranging from larger to smaller objects), and surface textures (from smooth surfaces to different complex textures). Some objects are also fragile (e.g. onion ring, egg, and cracker). All of these present challenges to the generalization and robustness of grasping. The position trajectory of dynamic transportation is shown in Fig. 13(a). We use this trajectory for evaluating all objects and methods. The only exception is the box containing beads, which undergoes a relative displacement along the Z-axis to accommodate its larger size. Objects are handed to the gripper by a person to facilitate grasping. Among evaluating for different methods, each object will be grasped with same configuration.
The task involves grasping different daily objects, each with different sizes, shapes, physical properties, and surface textures, and transporting them to another location. The task presents three main challenges. First, the gripper needs to apply an appropriate force to grasp the object without damaging it, but with enough force to hold it stably. Second, during the dynamic transportation of the object, the gripper needs to reactively adjust its behavior to maintain the object stably, particularly for heavier objects and objects with varying mass distribution. Third, the task requires generalizing feasible grasping to different objects with different physical properties, shapes, sizes, and surface textures.
@@
1) The detailed results of the LeTac-MPC implementation on the banana and box containing beads are shown in Fig. 7. When the object was not in contact with the gripper, the gripper width decreased due to the empty tactile feedback. When the object came into contact with the gripper, the gripper width gradually converged to a position based on the tactile feedback from the GelSight. After the controller converged, the manipulator dynamically transported the object to another location, and the tactile feedback changed due to rapid changes in the states the grasped object. Correspondingly, LeTac-MPC would enable the gripper reactively re-grasp the object to stabilize it in hand. Finally, after the object was transported to another location and become static, the gripper width converged.
@@
2) Fig. 8 shows the tactile images of all 10 daily objects in the experiments. Table IV shows the results and mean grasping forces of the experiments. Our proposed LeTac-MPC is able to generalize to different objects with different physical properties, sizes, shapes, and surface textures, as shown in Fig. 8. Items such as cracker (Fig. 8(d)), oion ring (Fig. 8(e)), and oreo (Fig. 8(f)) are smaller than the size of the gripper and also have complex textures and cluttered contours. The training dataset only includes four standardized blocks with different physical properties. This indicates the good generalization ability of LeTac-MPC. In Fig. 9, we show the gripper width curves of several experiments.
@@
3) As shown in Figs. 8(b) and 8(i), although the tactile feedback for soft objects like bagel and tomato is subtle, LeTac-MPC can still grasp robustly. However, both PD and MPC controllers rely on contact area as the feedback signal to compute control inputs, and therefore cannot function on soft objects that have subtle tactile feedback. This is because it is difficult to obtain a high-quality and stable contact area from the tactile feedback for soft objects. There are two popular ways to extract the contact area mask of the tactile image, thresholding depth images (Fig. 10) and difference images (Fig. 11). The depth image is the 3D reconstruction of the contact surface and can be obtained by the method in [49]. The difference image is obtained by subtracting the current frame image from the first frame image, with marker removal achieved by interpolation at the marker locations on both the current and first frames.
@@
We collect multiple tactile images for different objects in the experiment and present a sequence of results obtained by thresholding depth images and difference images by different values. As seen in Figs. 10 and 11, the same thresholding value cannot generalize to different objects with different physical properties. When the value is small, we cannot obtain a reasonable contact area for deformable objects. However, when the value is large, the contact area of rigid objects become larger than the reasonable area and has more noises. Furthermore, for bagel, regardless of how we adjust the value, the contact area always appears poor for computing control input. This is because the bagel is very soft, resulting in very subtle tactile feedback. The raw images are from MPC experimental records. This explains the results in Table IV. For PD, since it utilizes the same feedback signals and same control objectives as MPC, this can also explain the experimental results of PD.
In our implementation, we use the contact area from depth images as the feedback signal for MPC and PD. Additionally, 0.05 is the threshold we use. As shown in Fig. 10, 0.2 is the best threshold for rigid objects with textures and without textures.
@@
The poor contact area signal leads to poor grasping behavior. As shown in Fig. 9, MPC and PD for soft object grasping continually decreased the gripper width until it reached the force limit of the WSG 50-110 gripper, as the large force applied by the gripper cannot create a sufficient contact area mask. However, LeTac-MPC performs exceptionally well in grasping deformable objects, even when the tactile features are subtle.
@@
Note that for tomato grasping in Fig. 9, MPC succeeded while PD did not. This does not necessarily mean that MPC outperformed PD in grasping this object. As shown in Fig. 10, the contact area in the MPC experiment is also very unsatisfactory, which results in MPC convergence but with a large gripping force applied. Both PD and MPC calculate control inputs based on the contact area, which makes them both unsuitable for grasping deformable objects. The difference in experimental results for grasping tomatoes can be attributed to other minor factors, such as the grasping position of the tomato.
@@
4) Table IV shows that LeTac-MPC implements the smallest grasping force but performs the best on the grasping and transportation task. However, PD and MPC always implement very large force when grasping deformable objects, even when PD and MPC converge. The tomato case with MPC in Fig. 9 provides a good example. For open-loop grasping with large grasping force, it still fails to grasp the box containing beads stably during dynamic transportation. This is because the simple open-loop policy cannot re-grasp properly when the grasped object’s state changes suddenly. We can see the tactile image sequence of dropping the box in Fig. 18. Moreover, by comparing Figs. 8 and 18, we can see from the tactile feedback that open-loop grasping implements larger force than LeTac-MPC.
@@
5) In our implementation, due to the slow convergence rate of PD and MPC, we set the gripper to contract at a constant speed of 2.5 mm/s when there is no contact with the object. When the gripper is in contact with the object, we switch to the PD and MPC controllers. As shown in Fig. 9, the gripper width curves generated by PD and MPC both have a turning point (green circle), which represents the gripper width at the point of contact with the object. We can observe that after the turning point, PD and MPC gradually converge, but their convergence rate is slow. For LeTac-MPC, we do not set a constant contraction speed when the gripper is not in contact with an object. Instead, we apply LeTac-MPC to generate gripper motion through the entire grasping process. This is because LeTac-MPC has a faster convergence rate than PD and MPC, no matter if the gripper is contact with the object, as shown in Fig. 9.
@@
6) We can observe that the fluctuation after convergence of PD and MPC is smaller than that of LeTac-MPC. This is because the output of the NN has some randomness, but the output of model-based control is more stable. However, as shown in Figs. 7 and 9, the randomness in the output of our proposed network model does not affect grasping performance.
@@ V-B Extreme Case: Grasping a Peeled Boiled Egg
In this section, we use LeTac-MPC to grasp a peeled boiled egg, which is extremely soft and delicate. In fact, this experiment is not successful, but it shows a extreme case of LeTac-MPC.
The experimental visualizations and gripper width curves are shown in Fig. 12. We can see that, during the entire grasping process, the changes in tactile images are almost imperceptible to the human eye. This is because comparing with the elastomer of GelSight, the peeled boiled egg is much more softer. So no matter how much grasping force applied to the peeled biled egg, it can not make enough deformation on the elastomer of GelSight.
When the elastomer touch the egg for the first time, there is some subtle features showing on the tactile image. Therefore, the speed at which the gripper width decreases begins to vary, and it appears that the controller is trending towards convergence (point B in Fig. 12). However, it’s important to note that the controller does not actually converge; instead, the gripper width decreases very slowly with a minimal slope (point B in Fig. 12). This is because, for grasping other daily objects in Fig. 6, when the tactile image has features, the LeTac-MPC will make the gripper width to slowly decrease and tend towards convergence. During this slow decrease, as the grasping force increases, the features of the tactile image become more pronounced, which leads LeTac-MPC to gradually converge. However, as we mentioned, since a peeled boiled egg is too soft, no matter how much force is applied to it, its tactile features do not become stronger. Therefore, the gripper width continues to decrease at a very slow rate, and LeTac-MPC never converges. This ultimately leads to the egg being damaged (points C and D in Fig. 12).

### v2

@@ IV-B Model Predictive Control
TABLE IV: Results of daily objects grasping and transportation experiment. If an experiment fails, we indicate the corresponding table position with red font to highlight the failed result. If the experiment succeeds, we report the average value and standard deviation of the force during the experimental process after the grasping stabilizes. The bolded font represents the minimum grasping force for the same object. Here, the term “egg” refers to an egg with its shell, which is actually rigid. In the paper, we use “peeled boiled egg” to denote an egg without its shell.
@@
Figure 13: Position trajectories for the dynamic transportation (Section V-A), dynamic shaking (Section V-C), and obstacle collision (Section V-D) experiments. The maximum velocity of the trajectory for dynamic transportation is 1.40 m/s, with a maximum acceleration of 3.82 m/$\text{s}^{2}$. For dynamic shaking, the trajectory has a maximum velocity of 1.38 m/s and a maximum acceleration of 8.15 m/$\text{s}^{2}$. Lastly, the trajectory for obstacle collision has a maximum velocity of 0.47 m/s and a maximum acceleration of 0.65 m/$\text{s}^{2}$. This figure serves as a further supplement to the experiments, highlighting the dynamic nature and challenges of the trajectories involved.
@@
Figure 15: Gripper width curves of dynamic shaking experiments with MPC and PD control. We know that the gripper width and grasping force are negatively correlated (smaller gripper width means tighter grasping and larger grasping force). We can see from this figure that at the beginning of the experiments, the contact areas are too small compared to the raw images. Therefore, for MPC and PD, the initial converged grasping forces are actually larger than the proper grasping force.
@@
where $K_{c}$ is a scalar factor. Equation (48) assumes a linear relationship between the contact area $c$ and the gripper width $p$, which is a simple and valid approximation of complex gel dynamics. The linear model is a local approximation for a very short duration. In MPC, the model is not used for long-term predictions, as this would lead to significant errors. Our proposed MPC baseline method runs at a high frequency with receding horizon control, ensuring that we can continuously iterate with the local model to minimize model errors. Combining equations (5) and (48), we get the following model:
@@
$J(\mathbf{y}_{n},\mathbf{a}_{n})$ $=P\mathbf{e}_{n+N}^{T}{\mathbf{Q}}\mathbf{e}_{n+N}+\sum_{k=n}^{n+N-1}\mathbf{e}_{k}^{T}{\mathbf{Q}}\mathbf{e}_{k}+Q_{a}a_{k}^{2},$ (62)
@@
The optimization problem (68) is a QP. Unlike the differentiable MPC layer, we do not need to do backpropagation for the optimization problem (68). Here, we only need to solve the optimization problem (68) to compute the control inputs. Therefore, we use OSQP [50], an efficient QP solver to solve the optimization problem (68) in real-time. The parameters we use for MPC are shown in Tables II and III.
@@ IV-C Open-loop Grasping
For open-loop grasping, we use the force feedback of the WSG 50-110 gripper as the signal to choose the gripper width. We select 10 N as the threshold value. At the beginning of the grasping, the gripper width decreases, and once the force feedback is greater than 10 N, the gripper stops. Then the gripper width will remain in this position during the period of performing the task.
@@
It is important to note that even when we use 10 N as the static condition for the fingers in open-loop grasping, it does not imply that open-loop grasping can maintain a constant grasping force of 10 N. This variability is in part due to the inherent noise in the force feedback of the WSG gripper. Furthermore, when the force feedback reaches 10 N, the fingers are in motion, and the force exerted differs once they stop. Additionally, the final resting force varies with the shape and properties of the grasped object, as shown in Table IV.
@@ V Experiments
We conduct experimental validation of proposed LeTac-MPC on a WSG 50-110 parallel gripper and a Franka Panda manipulator, as shown in Fig. LABEL:fig:firstPage. We run LeTac-MPC at a frequency of 25 Hz. Using the official WSG 50 ROS package [^1] https://github.com/nalt/wsg50-ros-pkg, the maximum rate of the gripper is 30 Hz. Therefore, the gripper can track the motion generated by LeTac-MPC with reasonable saturation constraints. We perform three tasks, namely daily objects grasping and transportation (Section V-A), dynamic shaking (Section V-C), and obstacle collision (Section V-D), and compare LeTac-MPC with PD control, MPC, and open-loop grasping. To capture the tactile feedback, we mount a GelSight on one of the fingers of the WSG 50-110 gripper. A video includes these three tasks and comparisons with the baseline methods can be found in the supplemental material.
The WSG gripper we use has a built-in low-level controller, which can track position reference. Our designed reactive grasping controllers are high-level controllers. These high-level controllers produce reference motions for the gripper fingers with reasonable saturation constraints, which are then accurately tracked by the low-level controller. In the case of LeTac-MPC and MPC, the solution to the optimization problem is acceleration $a_{n}$, but our state also includes gripper width $p_{n}$ and velocity $v_{n}$. Eventually, we use $p_{n}$ from the state as the reference sent to the low-level controller for tracking. Although we do not directly track $v_{n}$ and $a_{n}$, since $p$ is derived from the integration of $v$ and $a$, what the low-level controller effectively tracks is the entire trajectory of the finger motion. As for the PD controller, at each moment, we calculate $v_{n}$ through feedback, which, after integration, gives us the position reference $p_{n}$ to be tracked by the low-level controller. Hence, the low-level controller actually tracks both $p$ and $v$.
@@ V-A Daily Objects Grasping and Transportation
We evaluate our proposed LeTac-MPC method and compare it with three baseline methods for a daily objects grasping and transportation task. As shown in Fig. 6, we select 10 daily objects. These objects differ in physical properties (such as stiffness, total weight, mass distribution, and coefficient of friction), sizes and shapes (ranging from larger to smaller objects), and surface textures (from smooth surfaces to different complex textures). Some objects are also fragile (e.g. onion ring, egg, and cracker). All of these present challenges to the generalization and robustness of grasping. The position of dynamic transportation is shown in Fig. 13(a). We use this trajectory for evaluating all objects and methods. The only exception is the box containing beads, which undergoes a relative displacement along the Z-axis to accommodate its larger size. Objects are handed to the gripper by someone to facilitate grasping. Among evaluating for different methods, each object will be grasped with the same configuration.
The task involves grasping different daily objects, each with different sizes, shapes, physical properties, and surface textures, and transporting them to another location. The task presents three main challenges. First, the gripper must apply the appropriate force to grasp the object without damaging it but with enough force to hold it stably. Second, during the dynamic transportation of the object, the gripper needs to reactively adjust its behavior to maintain the object stably, particularly for heavier objects and objects with varying mass distribution. Third, the task requires generalizing feasible grasping to different objects with different physical properties, shapes, sizes, and surface textures.
@@
1) The detailed results of the implementation of LeTac-MPC in banana and box containing beads are shown in Fig. 7. When the object was not in contact with the gripper, the gripper width decreased due to the empty tactile feedback. When the object came into contact with the gripper, the gripper width gradually converged to a position based on the tactile feedback from the GelSight. After the controller converged, the manipulator dynamically transported the object to another location, and the tactile feedback changed due to rapid changes in the states the grasped object. Correspondingly, LeTac-MPC would enable the gripper to reactively re-grasp the object to stabilize it in hand. Finally, after the object was transported to another location and became static, the gripper width converged.
@@
2) Fig. 8 shows the tactile images of all 10 daily objects in the experiments. Table IV shows the results and mean grasping forces of the experiments. Our proposed LeTac-MPC is able to generalize to different objects with different physical properties, sizes, shapes, and surface textures, as shown in Fig. 8. Items such as cracker (Fig. 8(d)), onion ring (Fig. 8(e)), and oreo (Fig. 8(f)) are smaller than the size of the gripper and also have complex textures and cluttered contours. The training dataset only includes four standardized blocks with different physical properties. This indicates a good generalizability of LeTac-MPC. In Fig. 9, we show the gripper width curves of several experiments.
@@
3) As shown in Figs. 8(b) and 8(i), although the tactile feedback for soft objects such as bagel and tomato is subtle, LeTac-MPC can still grasp robustly. However, both PD and MPC controllers rely on the contact area as the feedback signal to compute control inputs, and therefore cannot function on soft objects that have subtle tactile feedback. This is because it is difficult to obtain a high-quality and stable contact area from tactile feedback for soft objects. There are two popular ways to extract the contact area mask of the tactile image, thresholding depth images (Fig. 10) and difference images (Fig. 11). The depth image is the 3D reconstruction of the contact surface and can be obtained by the method in [51]. The difference image is obtained by subtracting the current frame image from the first frame image, with marker removal achieved by interpolation at the marker locations on both the current and the first frames.
@@
We collect multiple tactile images for different objects in the experiment and present a sequence of results obtained by thresholding depth images and difference images by different values. As seen in Figs. 10 and 11, the same thresholding value cannot be generalized to different objects with different physical properties. When the value is small, we cannot obtain a reasonable contact area for deformable objects. However, when the value is large, the contact area of rigid objects becomes larger than the reasonable area and has more noises. Furthermore, for bagel, regardless of how we adjust the value, the contact area always appears poor for computing control input. This is because the bagel is very soft, resulting in very subtle tactile feedback. The raw images are from MPC experimental records. This explains the results in Table IV. For PD, since it utilizes the same feedback signals and control objectives as MPC, this can also explain the experimental results of PD.
In our implementation, we use the contact area from depth images as the feedback signal for MPC and PD. Additionally, 0.05 is the threshold that we use. As shown in Fig. 10, 0.05 is the optimal threshold for rigid objects with textures and without textures based on our experiments in Fig. 10. In Fig. 10, we can see that using 0.025 as the threshold introduces noise in non-contact areas for hard objects such as cracker and avocado (visible on the right side), and the contact area for cracker becomes elongated, which does not match the actual ground truth. Although tomato shows improved contact area quality at 0.025, the quality for bagel remains poor. On the other hand, a 0.05 threshold performs better for rigid objects (cracker and avocado). Therefore, we chose 0.05, which generally works for rigid objects.
@@
The poor contact area signal leads to poor grasping behavior. As shown in Fig. 9, MPC and PD for soft object grasping continuously decreased the gripper width until it reached the force limit of the WSG 50-110 gripper, as the large force applied by the gripper cannot create a sufficient contact area mask. However, LeTac-MPC performs exceptionally well in grasping deformable objects, even when tactile features are subtle.
@@
Note that for tomato grasping in Fig. 9, MPC succeeded while PD did not. This does not necessarily mean that MPC outperformed PD in grasping this object. As shown in Fig. 10, the contact area in the MPC experiment is also very unsatisfactory, which results in MPC convergence, but with a large gripping force applied. Both PD and MPC calculate control inputs based on the contact area, which makes both unsuitable for grasping deformable objects. The difference in the experimental results for grasping tomato can be attributed to other minor factors, such as the grasping position of the tomato.
@@
4) Table IV shows that LeTac-MPC demonstrates statistically smaller grasping forces for most objects and performs the optimal on the grasping and transportation task. For very soft objects, such as tomato and bagel, PD and MPC tend to apply excessive grasping forces. Often, these forces reach the gripper’s force limit. For open-loop grasping with large grasping force, it still fails to grasp the box containing beads stably during dynamic transportation. This is because the simple open-loop policy cannot re-grasp properly when the grasped object’s state changes suddenly. We can see the tactile image sequence of dropping the box in Fig. 18. Moreover, for certain objects in which the force applied during open-loop grasping is much greater than that in LeTac-MPC, we can see differences between tactile images Figs. 8 and 18. For example, by comparing Figs. 8 and 18, it is evident that the tactile image of the egg in Fig. 18 has a larger contact area and clearer features, indicating that the force applied in open-loop grasping is greater.
@@
5) In our implementation, due to the slow convergence rate of PD and MPC, we set the gripper to contract at a constant speed of 2.5 mm/s when there is no contact with the object. When the gripper is in contact with the object, we switch to the PD and MPC controllers. As shown in Fig. 9, the gripper width curves generated by PD and MPC both have a turning point (green circle), which represents the gripper width at the point of contact with the object. We can observe that after the turning point, PD and MPC gradually converge, but their convergence rate is slow. For LeTac-MPC, we do not set a constant contraction speed when the gripper is not in contact with an object. Instead, we apply LeTac-MPC to generate gripper motion throughout the grasping process. This is because LeTac-MPC has a faster convergence rate than PD and MPC, no matter if the gripper is in contact with the object, as shown in Fig. 9.
@@
6) We can observe that the fluctuation after convergence of PD and MPC is smaller than that of LeTac-MPC. This is because the output of the NN has some randomness, but the output of the model-based controllers is more stable. However, as shown in Figs. 7 and 9, the randomness in the output of our proposed network model does not affect grasping performance.
@@ V-B Extreme Case: Grasping a Peeled Boiled Egg
In this section, we use LeTac-MPC to grasp a peeled boiled egg, which is extremely soft and delicate. In fact, this experiment is not successful, but shows an extreme case of LeTac-MPC.
The experimental visualizations and gripper width curves are shown in Fig. 12. We can see that, during the entire grasping process, the changes in tactile images are almost imperceptible to the human eye. This is because comparing with the elastomer of GelSight, the peeled boiled egg is much more softer. So, no matter how much grasping force is applied to the peeled biled egg, it cannot make enough deformation on the elastomer of GelSight.
When the elastomer touches the egg for the first time, there are some subtle features that show up in the tactile image. Therefore, the speed at which the gripper width decreases begins to vary, and it appears that the controller is trending toward convergence (point B in Fig. 12). However, it is important to note that the controller does not actually converge; instead, the gripper width decreases very slowly with a minimal slope (point B in Fig. 12). This is because, for grasping other daily objects in Fig. 6, when the tactile image has features, LeTac-MPC will slowly decrease the gripper width and tend toward convergence. During this slow decrease, as the grasping force increases, the features of the tactile image become more pronounced, leading LeTac-MPC to gradually converge. However, as we mentioned, since a peeled boiled egg is too soft, no matter how much force is applied to it, its tactile features do not become stronger. Therefore, the gripper width continues to decrease at a very slow rate, and LeTac-MPC never converges. This ultimately leads to the egg being damaged (points C and D in Fig. 12).